A Ferris wheel undergoes a constant angular acceleration as it starts from rest and builds up to a final angular speed of 1.9 rad/s while rotating through an angular displacement of 4.9 rad. Find its angular acceleration.
Answer #1

w0 = 0 and wf = 1.9 rad/s

and theta = 4.9 rad  

Applying wf^2 - wi^2 = 2 alpha theta

1.9^2 - 0^2 = 2(4.9) (alpha)

alpha = 0.368 rad/s^2 ..........Ans